# How Exclusive are Ethereum Transactions?

### v2

@@ Abstract
We analyze 15,097 blocks proposed for inclusion in Ethereum’s blockchain over an eight-minute window on December 3, 2024, during which 38 blocks were added to the chain. We classify transactions as exclusive—appearing only in blocks from a single builder—or private — absent from the public mempool but included in blocks from multiple builders. We find that exclusive transactions account for between 77.2% and 84% of the total fees paid by transactions in winning blocks. Moreover, we show that exclusivity cannot be fully attributed to persistent relationships between senders and builders: only about 7% of all on-chain exclusive transaction value originates from senders who route exclusively to one builder. Finally, we observe that transaction exclusivity is dynamic. Some transactions are exclusive at the start of a bidding cycle but later appear in blocks from multiple builders. Other transactions remain exclusive to a losing builder for two or three cycles before appearing in the public mempool. These transactions are therefore delayed and then exposed to potential attacks.
@@ 1 Introduction
This system, known as Proposer-Builder Separation (PBS), has been effective at selecting the most valuable transactions for inclusion on-chain. Its downside, however, is the extreme concentration in the builders’ market: the top builder (Titan) produces about 50% of blocks, while the top three account for more than 90%. [^1] For up-to-date statistics, see https://www.relayscan.io/. Note that Buildernet is a joint project by Flashbots (who operated their own independent builder until mid-2025) and Beaverbuild (who both participate in Buildernet and run their own builder). This concentration gives a few entities disproportionate influence over which transactions are included, threatening Ethereum’s foundational goals of decentralization and permissionlessness.
The leading explanation for why PBS produced such concentration rests on exclusive transactions and economies of scale: the more often a builder wins, the more attractive it becomes as a destination for order flow that is not shared with competitors, further increasing its wins in future auctions. Quantifying the importance of exclusive transactions, however, is difficult because on-chain data alone cannot reveal whether a transaction was privately shared with multiple builders or only with the eventual winner.
Our paper addresses this challenge using a novel dataset containing 15,097 blocks proposed for inclusion in Ethereum during an eight-minute window on December 3, 2024, corresponding to 39 new blocks. [^2] The dataset includes all blocks (winning and non-winning) submitted via the “Agnostic relay” during the study period. They are therefore not a random sample of all proposed blocks (see Section 3 for details and limitations). To our knowledge, together with the companion paper Canidio and Pahari (2025), this is the first study to analyze the content of non-winning Ethereum blocks. The dataset contains 10,793 unique transactions (identified by their hash) and 2,380,014 transaction–block pairings. By comparing the content of proposed blocks with that of the public mempool (sourced from https://mempool.guru/), we classify each transaction as:
  • Public: it appeared in the public mempool;
  • Private: it appeared in blocks from multiple builders but not in the public mempool;
  • Exclusive: it appeared only in blocks from a single builder.
We apply this classification both globally and within shorter time intervals. Applying the classification over shorter intervals allows us to track transitions between statuses over time. For instance, a transaction may be exclusive initially but later become public or private.
@@
  • functionally identical to each other, executing the same operation with the same fee;
@@
While this log-based analysis refines our classification of exclusivity, the number of affected transactions is small and does not alter our main conclusion: even if all such transactions were reclassified from “exclusive” to “private,” exclusive transactions would still account for a large share (77.2%) of total fees paid by transactions included in winning blocks.
@@
To investigate the origin of exclusive transactions, we analyze the relationship between transaction senders and builders. Counting transactions alone, we find that the vast majority of exclusive transactions can be attributed to senders that interact exclusively with a single builder. However, when we shift focus from transaction counts to transaction value and consider only those that are ultimately included on-chain, the picture changes. Under our preferred operational definition of a “transaction sender,” only about 7% of the total value of exclusive transactions can be attributed to exclusive relationships between searchers and builders. Such relationships therefore appear to play a limited role in explaining the overall value of exclusive transactions.
@@
Finally, we study whether transactions change status over time, starting with changes across bidding cycles. Among transactions that first appear as exclusive, 162 remain exclusive over multiple cycles, 174 transition to private in a subsequent cycle, and 12 later appear in the public mempool. These 12 transactions appear to be user-generated, as they involve swaps on DEX pools via well-known routers such as 0x and 1inch. Their initial exclusivity is puzzling, as ordinary users typically gain no benefit from sending transactions directly to a single builder instead of submitting them to a private relay or the public mempool. Even more surprisingly, these transactions eventually surface in the public mempool, indicating that they were initially delayed and later exposed to potential attacks (one was in fact the victim of a sandwich attack). In four cases, the same builder who initially had exclusive access ultimately included the transaction on-chain. We speculate that a private mempool operator may have shared these transactions exclusively with one builder, reverting to the public mempool when that builder failed to win. [^3] Private mempools are services that allow users to send transactions directly to one or more builders without broadcasting them to the public mempool. At present, there are five major private mempool operators. Four of them—MEV Blocker, Flashbots Protect, Merkel, and Blink—state in their documentation that they share transactions privately with multiple builders (on the differences between those operators, see Janicot and Vinyas, 2025). Instead, the fifth operator, Metamask Smart Transaction, does not disclose its sharing policy.
@@
We also observe transactions changing status within a single bidding cycle. Specifically, we identify 19 cases in which transactions are initially exclusive to one builder at the start of the auction cycle but later, within the same cycle, appear in blocks proposed by other builders. Most of these transactions were submitted via Flashbots Protect: they were initially exclusive to the Flashbots builder and only later propagated more broadly. We speculate that the Flashbots builder may temporarily retain transactions submitted through Flashbots Protect as exclusive, releasing them to other builders once it becomes apparent that it is unlikely to win the block.
@@
To our knowledge, Yang et al. (2024) is the only other paper that examines the content of non-winning blocks. They study builder competition by comparing bids with block revenues (i.e., total payments received by builders), but do not analyze the underlying transactions.
Other strands of the literature explore builders’ competition from different angles. Wu et al. (2024a, b) analyze bidding strategies, both theoretically and via simulation. Titan and Frontier Research (2023); Öz et al. (2024) show that builder profits hinge on how they source transactions, using a classification approach related to ours. Gupta et al. (2023) formalizes how exclusive transactions can drive concentration in the equilibrium of the builders’ market. Finally, Bahrani et al. (2024) model heterogeneous builders and derive conditions under which both builder and proposer markets become concentrated.
@@ 4 Results
(a) 11 seconds
(b) 12 seconds
(c) 13 seconds
(d) 14 seconds
Figure 1: Proportion of transactions shared between blocks by different builders, at different moments in the bidding cycle.
As a first step, Figure 1 shows that blocks proposed by different builders contain different transactions. To study these differences, we label transactions using MempoolGuru as either public (it appeared in the public mempool) or non-public (it appeared in a block but not in the mempool). We further split non-public transactions into private (present in blocks from multiple builders) and exclusive (present only in blocks from a single builder).
@@ 4.2 Relationship between searchers and builders
5
6
(1.22%)
19
(3.9%)
6
(3.9%)
19
(6.0%)
1.14
(9.3%)
1.28
(10.4%)
1.14
(13.87%)
1.28
(10.47%)
Table 2: Exclusive transactions by sender, grouped by how many distinct builders each sender interacts with. Columns distinguish senders defined by from vs. to addresses, with and without a minimum threshold ($\geq 5$ transactions). The left panel reports the count of exclusive transactions that landed on-chain, with the share of total exclusive transactions included on-chain underneath. The right panel reports total fees in ETH among the transactions that landed on-chain, with the share of total underneath.
Tables 1 and 2 group transactions by the number of builders each sender interacts with, using both the to and the from definitions of “sender” (see the previous subsection) and with and without filtering out senders who send fewer than five transactions. [^10] The rationale behind the filtering is that a sender with only one or two exclusive transactions will mechanically appear to work exclusively with a single builder, which may reflect limited data rather than a genuine relationship. Table 1 reports the resulting transaction counts and the total fees paid by those transactions, while Table 2 reports the same but only for transaction that are ultimately included on chain.
@@
In Table 1 transaction counts vary somewhat with the definition of “sender” but the fee data are more robust: by value, roughly half of all exclusive transactions arise from exclusive relationships between senders and builders. However, once we restrict to on-chain transactions, this share drops sharply. Table 2 shows that if we use to addresses—which better identify searchers—then only around 7% of the value of exclusive transactions that land on-chain can be traced to exclusive relationships. Taken together, these results suggest that exclusive relationships between senders and builders play only a limited role in explaining why, as noted earlier, 85% of the value of on-chain transactions comes from exclusive transactions.
@@ 5 Conclusion
Together with its companion paper Canidio and Pahari (2025), this study provides the first systematic analysis of non-winning Ethereum blocks, enabling us to distinguish between public, private, and exclusive transactions. We show that although exclusive transactions represent a minority of all transactions included on chain, they account for the vast majority of builders’ revenues—between 77.2% and 84% of total fees in winning blocks, depending on the methodology. While exclusive relationships between senders and builders are common, they explain only a small share (approximately 7%) of the total value of exclusive transactions included on-chain. By demonstrating that exclusive transactions are responsible for the vast majority builder revenues, our results support the view that exclusivity is a key driver of market concentration and helps explain the dominance of a few builders.

### v3

@@ Abstract
We analyze 15,097 blocks proposed for inclusion in Ethereum’s blockchain over an eight-minute window on December 3, 2024, during which 38 blocks were added to the chain. We classify transactions as exclusive—appearing only in blocks from a single builder—or private — absent from the public mempool but included in blocks from multiple builders. We find that, depending on the methodology, exclusive transactions account for between 77.2% and 84% of the total fees paid by transactions in winning blocks. Moreover, we show that exclusivity cannot be fully attributed to persistent relationships between senders and builders: only between 7% and 8.4% of all on-chain exclusive transaction value originates from senders who route exclusively to one builder. Finally, we observe that transaction exclusivity is dynamic. Some transactions are exclusive at the start of a bidding cycle but later appear in blocks from multiple builders. Other transactions remain exclusive to a losing builder for two or three cycles before appearing in the public mempool. These transactions are therefore delayed and then exposed to potential attacks.
@@ 1 Introduction
This system, known as Proposer-Builder Separation (PBS), has proven effective at selecting the most valuable transactions for inclusion on chain. Its main drawback, however, is the resulting concentration in the builders’ market: at present, the top builder (Titan) produces about 50% of all blocks, while the top three account for over 90%. [^1] For up-to-date statistics, see https://www.relayscan.io/. Buildernet is a joint project by Flashbots (which operated its own independent builder until mid-2025) and Beaverbuild (which both participates in Buildernet and runs its own builder). This concentration grants a few entities disproportionate influence over which transactions are included, undermining Ethereum’s foundational principles of decentralization and permissionlessness.
The prevailing explanation for the concentration observed under PBS centers on exclusive transactions, i.e., those available to a single builder only. Transactions shared among multiple builders should not affect the equilibrium outcome of the PBS auction, since all builders should raise their proposed payments to the proposer by an amount equal to these transactions’ fees, which ultimately accrue to the winning builder. As a result, the auction outcome is primarily determined by exclusive transactions. Furthermore, the more frequently a builder wins, the more attractive it becomes as a destination for non-shared order flow, further increasing its likelihood of winning future auctions and leading to concentration in the builders’ market (see Kilbourn, 2022).
Despite their central role in the PBS mechanism, quantifying the importance of exclusive transactions has proved difficult, since on-chain data alone cannot reveal whether a transaction was privately shared among multiple builders or only with the winner. Our paper takes a first step toward addressing this challenge using a novel dataset comprising 15,097 blocks proposed for inclusion in Ethereum over an eight-minute window on December 3, 2024, during which 39 new blocks were added. [^2] The dataset includes all blocks (winning and non-winning) submitted via the “Agnostic Relay” during the study period. They are therefore not a random sample of all proposed blocks (see Section 3 for details and limitations). To our knowledge, together with the companion paper Canidio and Pahari (2025), this is the first study to analyze the content of non-winning Ethereum blocks. The dataset contains 10,793 unique transactions (identified by hash) and 2,380,014 transaction–block pairings. By comparing the content of proposed blocks with the public mempool (sourced from https://mempool.guru/), we classify each transaction as:
  • Public: appeared in the public mempool;
  • Private: appeared in blocks from multiple builders but not in the public mempool;
  • Exclusive: appeared only in blocks from a single builder.
We apply this classification both globally and within shorter time intervals, allowing us to track how transaction statuses evolve. For example, a transaction may initially be exclusive but later become public or private.
@@
  • functionally identical to each other, executing the same operation and paying the same fee;
@@
While this log-based analysis refines our classification of exclusivity, the number of affected transactions is small and does not alter our main conclusion: even if all such transactions were reclassified from “exclusive” to “private,” exclusive transactions would still account 77.2% of total fees paid by transactions included in winning blocks.
@@
To investigate the origin of exclusive transactions, we then analyze the relationship between transaction senders and builders. Counting transactions alone, we find that the vast majority of exclusive transactions can be attributed to senders that interact exclusively with a single builder. However, when we shift focus from transaction counts to transaction value and consider only those that are ultimately included on-chain, the picture changes. Under our preferred operational definition of a “transaction sender,” only about 7% of the total value of exclusive transactions can be attributed to exclusive relationships between searchers and builders. Such relationships therefore appear to play a limited role in explaining the overall value of exclusive transactions.
@@
Finally, we study whether transactions change status over time, starting with changes across bidding cycles. Among transactions that first appear as exclusive, 162 remain exclusive over multiple cycles, 174 transition to private in a subsequent cycle, and 12 later appear in the public mempool. These 12 transactions appear to be user-generated, as they involve swaps on DEX pools via well-known routers such as 0x and 1inch. Their initial exclusivity is puzzling, as ordinary users typically gain no benefit from sending transactions to a single builder instead of submitting them to a privately to multiple builders or to the public mempool. Even more surprisingly, these transactions eventually appear in the public mempool, indicating that they were initially delayed and later exposed to potential attacks (one was in fact the victim of a sandwich attack). In four cases, the same builder who initially had exclusive access ultimately included the transaction on-chain. We speculate that a private mempool operator may have shared these transactions exclusively with one builder, reverting to the public mempool when that builder failed to win. [^3] Private mempools are services that allow users to send transactions directly to one or more builders without broadcasting them to the public mempool. During the study period, there were five major private mempool operators. Four of them — MEV Blocker, Flashbots Protect, Merkel, and Blink — state in their documentation that they share transactions privately with multiple builders (on the differences between those operators, see Janicot and Vinyas, 2025). Instead, the fifth operator, Metamask Smart Transaction, does not disclose its sharing policy.
@@
We also observe transactions changing status within a single bidding cycle. Specifically, we identify 19 cases in which transactions are initially exclusive to one builder at the start of the auction cycle but later, within the same cycle, appear in blocks proposed by other builders. Most of these transactions were submitted via Flashbots Protect: they were initially exclusive to the Flashbots builder and only later propagated more broadly. Our interpretation is that Flashbots builder temporarily retains transactions submitted through Flashbots Protect as exclusive, releasing them to other builders once it becomes apparent that it is unlikely to win the block.
@@
To our knowledge, Yang et al. (2025) is the only other paper that examines the content of non-winning blocks. They study builder competition by comparing bids with realized block revenues (i.e., total payments received by builders), but do not analyze individual transactions. The authors also propose a heuristic to infer exclusive transactions from bid-level data. In their approach, exclusive transactions are defined as private transactions included in a winning block from an originator who is pivotal—that is, if the transactions from that originator were removed, the block’s value would fall below the second-highest builder’s bid (multiplied by a calibration threshold). This method effectively identifies some sources of exclusive order flow but has two main limitations. First, a set of transactions may appear pivotal yet be present also in the second-most valuable block. Second, a truly exclusive transaction may fail to appear pivotal if the runner-up block contains a different exclusive transaction of comparable value.
Other studies attempting to identify “private” and “exclusive” transactions have been limited by the difficulty of disentangling them using on-chain data alone. For example, Thiery (2023) labels as exclusive any transaction included on chain without first appearing in the public mempool—what we would classify as either private or exclusive—and reports that such transactions account for roughly 80% of the total value paid to builders, consistent with our findings (a similar estimate is reported by Yang et al., 2025). Titan and Frontier Research (2023) studies flows from identified searchers (bots) and documents a positive relationship between builder market share and flow from searchers, without distinguishing between private and exclusive transactions. Finally, Öz et al. (2024) shows that builders’ profitability correlates with access to exclusive transactions. However, their definition of “exclusive” includes all transactions not seen in the public mempool and not identified shared via MEV Blocker or Flashbots protect, two private mempool services. In our dataset, these two sources account for 320 of the 810 transactions that we classify as private.
A growing theoretical literature examines builder competition and the role of exclusive transactions in shaping the outcome of the PBS auction. In particular, Gupta et al. (2023) and Capponi et al. (2024) formalize how access to exclusive transactions can lead to concentration in the builder market. Bahrani et al. (2024) introduce heterogeneity among builders and derive conditions under which both builder and proposer markets become concentrated. Wu et al. (2024a, b) analyze builders’ bidding strategies, combining theoretical modeling with simulation-based validation. Lehar and Parlour (2023) and Capponi et al. (2025) study the tradeoff between submitting transactions via the public mempool versus private channels, and characterize the resulting equilibrium sorting of transactions; Capponi et al. (2025) also provides empirical support for its model using data from Flashbots Protect. Similarly, Azar et al. (2025) develop a theory in which the choice between public and exclusive submission depends on the exogenous value of private information; when this value is sufficiently high, the model predicts concentration in the builder market. They test this theory empirically by classifying transactions as “public” or “private,” under the assumption that private transactions are available exclusively to a single builder.
@@ 4 Results
(a) Percentage of Transaction
(b) Percentage of Fees
Figure 1: Exclusive, non-exclusive (public and private), and non-identified transactions for each winning block. For each block, we report both the share of transactions and the share of total fees attributable to each category. Blocks are numbered 21,322,622 through 21,322,660.
We label the transactions in our dataset using MempoolGuru as either public (it appeared in the public mempool) or non-public (it appeared in a block but not in the mempool). We further split non-public transactions into private (present in blocks from multiple builders) and exclusive (present only in blocks from a single builder). Note that transaction that appear in winning blocks during the study period but do not appear in our dataset (because included only in blocks not submitted to Agnostic relay) cannot be categorized.
@@
To further support our hypothesis, Figure 1 reports, for each winning block in our study period, the share of transactions that we classify as exclusive, as non-exclusive, or as not identifiable (because they do not appear in our dataset). Consistent with the discussion above, the only winning blocks with no exclusive transactions are the three built by Beaverbuild (21322626, 21322650, and 21322653). By contrast, the remaining 19 winning blocks—all produced by Titan or Rsync—almost never contain transactions that we cannot identify: 17 of them contain none, one block contains a single non-identified transaction of negligible value, and only one block (21322659 by Rsync) contains both economically meaningful exclusive transactions and non-identified transactions. Taken together, these patterns suggest a sharp separation across builders: Beaverbuild’s winning blocks contain non-identified transactions but no exclusive ones, whereas blocks built by the other major builders contain exclusive transactions but (almost) no non-identified ones. This supports our hypothesis that most transactions we cannot identify are exclusive to Beaverbuild. Finally, if we reclassify all transactions appearing in winning blocks but missing from our dataset as exclusive, the share of winning-block value attributable to exclusive transactions rises to 87%.
@@ 4.2 Relationship between searchers and builders
Table 2: Exclusive transactions by sender, grouped by how many distinct builders each sender interacts with. Columns distinguish senders defined by from vs. to addresses, with and without a minimum threshold ($\geq 5$ transactions). The left panel reports the count of exclusive transactions that included on chain, with the share of total exclusive transactions included on-chain underneath. The right panel reports total fees in ETH among the transactions included on chain, with the share of total underneath.
@@
In Table 1 transaction counts vary somewhat with the definition of “sender” but the fee data are more robust: by value, roughly half of all exclusive transactions arise from exclusive relationships between senders and builders. However, once we restrict to on-chain transactions, this share drops sharply. Table 2 shows that if we use to addresses—which better identify searchers—then only around 7% of the value of exclusive transactions that land on-chain can be traced to exclusive relationships. Hence, exclusive relationships between senders and builders play only a limited role in explaining why, as noted earlier, 85% of the value of on-chain transactions comes from exclusive transactions. [^11] Tables 4 and 5 in Appendix replicate Tables 1 and 2, but excluding from the set of exclusive transactions those that we previously identified as “partially exclusive” or “similar transactions”. The main result is unchanged: under this alternative categorization, only around 8.4% of the value of exclusive transactions that land on-chain can be traced to exclusive relationships.
@@ 5 Conclusion
Together with its companion paper Canidio and Pahari (2025), this study provides the first systematic analysis of non-winning Ethereum blocks, enabling us to distinguish between public, private, and exclusive transactions. We show that although exclusive transactions represent a minority of all transactions included on chain, they account for the vast majority of builders’ revenues—between 77.2% and 84% of total fees in winning blocks, depending on the methodology. While exclusive relationships between senders and builders are common, they explain only a small share (approximately 7%) of the total value of exclusive transactions included on-chain. By demonstrating that exclusive transactions are responsible for the vast majority builders’ revenues, our results support the view that exclusivity is a key driver of the PBS auction dynamics and the resulting dominance of a few builders.
@@
Appendix 0.A Additional tables
# of builders Transactions Fees (ETH)
From To From $\geq$5 To $\geq$5 From To From $\geq$5 To $\geq$5
1
3552
(91.22%)
3098
(79.56%)
2923
(82.05%)
2803
(100.0%)
88.7
(82.43%)
68.52
(63.68%)
64.36
(79.56%)
60.59
(61.7%)
2
225
(5.78%)
563
(14.46%)
159
(82.05%)
531
(100.0%)
11.32
(10.52%)
29.95
(27.83%)
8.97
(11.09%)
28.46
(28.98%)
3
91
(2.34%)
21
(0.54%)
87
(82.05%)
21
(100.0%)
5.5
(5.11%)
3.85
(3.58%)
5.48
(6.77%)
3.85
(3.92%)
4
16
(0.41%)
197
(5.06%)
16
(82.05%)
197
(100.0%)
1.26
(1.17%)
4.06
(3.77%)
1.26
(1.56%)
4.06
(4.13%)
Table 4: Exclusive transactions by sender, grouped by how many distinct builders each sender interacts with. Here we exclude from the set of exclusive transactions those we identified as similar or partially exclusive. Columns distinguish senders defined by from vs. to addresses, with and without a minimum threshold ($\geq 5$ transactions). The left panel reports transaction counts with the share of total transactions underneath. The right panel reports total fees in ETH with the share of total fees underneath.
# of builders Transactions Fees (ETH)
From To From $\geq$5 To $\geq$5 From To From $\geq$5 To $\geq$5
1
384
(93.89%)
334
(81.66%)
66
(23.47%)
165
(76.28%)
8.62
(76.13%)
0.95
(8.39%)
2.01
(41.75%)
0.87
(7.15%)
2
19
(4.65%)
66
(16.14%)
20
(23.47%)
125
(76.28%)
0.09
(0.79%)
7.74
(68.36%)
0.13
(2.7%)
8.53
(70.11%)
3
4
(0.98%)
3
(0.73%)
5
(23.47%)
3
(76.28%)
1.5
(13.25%)
1.49
(13.16%)
1.54
(31.99%)
1.49
(12.25%)
4
2
(0.49%)
6
(1.47%)
5
(23.47%)
19
(76.28%)
1.12
(9.89%)
1.14
(10.07%)
1.13
(23.47%)
1.28
(10.52%)
Table 5: Exclusive transactions by sender, grouped by how many distinct builders each sender interacts with. Here we exclude from the set of exclusive transactions those we identified as similar or partially exclusive. Columns distinguish senders defined by from vs. to addresses, with and without a minimum threshold ($\geq 5$ transactions). The left panel reports the count of exclusive transactions included on-chain, with the share of total exclusive transactions included on-chain underneath. The right panel reports total fees in ETH among the transactions included on-chain, with the share of total underneath.